OpenStudy (anonymous):

Solve, using the substitution method. 3x + 2y = 11 x 2 = 4y (Points : 1) The solution is (0, 2). The solution is . ( 12, -5/2) The solution is . (0, -1/2 The solution is . (4,-1/2)

OpenStudy (alfie):

What's that over there in the second equation, a plus?

OpenStudy (anonymous):

3x + 2y = 11 x - 2 = 4y srry everyone

OpenStudy (alfie):

Okay, first thing to do is to observe, you clearly see that the second equation is easier, so you isolate the x down there... x = 4y+2. Now plug this thing in the first one... 3(4y+2) +2y = 11 => 12y +6 +2y = 11 => 14y = 5 => y = 5/14 x = 24/7 Weird it doesn't match with any of the possible answers though, you sure you wrote it correctly?

OpenStudy (anonymous):

ya it is wierd to me too!